OpenStudy (anonymous):

The equation below represents Function A and the graph represents Function B: Function A f(x) = x - 9 Function B graph of line going through ordered pairs -1, -3 and 2, 3 Which equation best compares the slopes of the two functions? @.Sam.

OpenStudy (anonymous):

i will giv medal

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

r there more information?

OpenStudy (anonymous):

no

OpenStudy (jordanloveangel):

in order to answer this question, we will need to figure out what the slope is for function A and what the slope is for function B, looking at the equation for function A, its f(x) = x - 9 and compare it to the y-intrercept form that tells you where to look to find the slope of an equation: y = mx+b where the slope, m, is the number that is infront of x, can you tell me what the slope is of the function: f(x) = x - 9?

OpenStudy (anonymous):

The equation below represents Function A and the graph represents Function B: Function A f(x) = x - 9 Function B graph of line going through ordered pairs -1, -3 and 2, 3 Which equation best compares the slopes of the two functions? Slope of Function B = 2 × Slope of Function A. Sope of Function A = Slope of Function B Slope of Function A = 2 × Slope of Function B Slope of Function B = - Slope of Function A

OpenStudy (jordanloveangel):

Slope of Function A = 2 × Slope of Function B because fn a slope is 6and from the graph fn b slope is more than 2 so..... The slope of graph A, has a slope of 6, as y = 6x - 1 is slope of 6. The graph of B is a line with slope of 3, because for every x to the right, y goes up 3.

OpenStudy (jordanloveangel):

does this help

OpenStudy (anonymous):

i guess